OpenStudy (anonymous):

x^2-y^2=x+y+1. I know this is a hyperbola, but can u put it in the standard equation

OpenStudy (helder_edwin):

complete the squares.

OpenStudy (helder_edwin):

give me a second

OpenStudy (helder_edwin):

i got \[\large (x-1/2)^2-(y+1/2)^2=1 \]

OpenStudy (anonymous):

how did u get the one

OpenStudy (helder_edwin):

i will do all the steps \[\large x^2-y^2=x+y+1 \] \[\large x^2-x-y^2-y=1 \] \[\large (x^2-x+(1/2)^2)-(y^2+y+(1/2)^2)=1+(1/2)^2-(1/2)^2=1 \] \[\large (x-1/2)^2-(y+1/2)^2=1 \]

OpenStudy (anonymous):

thank u, I forgot to subtract the 1/4 so I ended up with 3/2 on the left. that is where I got stuck

OpenStudy (helder_edwin):

u r welcome

OpenStudy (anonymous):

I was right that it is a hyperbola

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

my friend and I out this ellipise in its standard form x^2+4y^2=2x+8y-1. she got this answer (x-1)^2 + 4(y-1)^2 =4. I got (x-1)^2/(16) + (y-4)^2/(4)= 1. which is correct @helder_edwin

OpenStudy (helder_edwin):

give a minute

OpenStudy (helder_edwin):

i got the first one: \[\large \frac{(x-1)^2}{4}+(y-1)^2=1 \]

OpenStudy (anonymous):

how

OpenStudy (helder_edwin):

\[\large x^2+4y^2=2x+8y-1 \] \[\large (x^2-2x+1)+4(y^2-2y+1)=-1+1+4=4 \] \[\large (x-1)^2+4(y-1)^2=4 \] \[\large \frac{(x-1)^2}{4}+(y-1)^2=1 \]

OpenStudy (anonymous):

compute limit (as t approaches 0) : sec^2t -1/ (t). I got 0, just want to confirm

OpenStudy (helder_edwin):

i got zero.

OpenStudy (anonymous):

compute limit (as t approaches 0) : sect -1/ (t). I got 0 too, just want to confirm

OpenStudy (anonymous):

I used the derivative. did u

OpenStudy (anonymous):

if not can u show me the other way u used

OpenStudy (helder_edwin):

i got zero. yes i used L'Hôpital's rule

OpenStudy (anonymous):

my teacher want us to find the derivative of this problem using the limit rule, but I get lose using the limit rule: (1/(x+1))